x= 1, 4, 5, 6, 8
y= _, 18, 23, 28, 38
I have to solve for y. I do not the equation to solve this problem. My book calls this problem graphing functions.

Answer by khwang(438)

This question is to find the linear function of y in terms of x.
Assume (y-y1)/(x-x1) = m , since when x=4, y =18 and when x=5,y=23.
slope m = (23-18)/(5-4) = 5.
We have (y-18)/(x-4) = 5, or y-18 = 5(x-4)
or y = f(x) = 5x -2.
Check: f(6) = 30-2 =28 = y(OK),
f(8) = 40-2 =38 (OK)